# Dichography: Two-frame Ultrafast Imaging from a Single Diffraction Pattern

### v2

@@ III.1 Seeing double from a single shot
The last example on “double-hit” data is reported in Fig. 3d, and presents many peculiarities of the samples shown before in one. The two frames of the reconstruction have a significantly different spatial extension, $80\text{\,}\mathrm{nm}$ for the smaller and up to a maximum of $200\text{\,}\mathrm{nm}$ for the larger. The smaller density has a simple cubic architecture, whereas the larger one can be interpreted as an agglomerate of four nanocubes. The shape of three cubes in the agglomerate is well imprinted into the overall outline, while the presence of the fourth cube is only revealed thanks to the successful reconstruction of the density values inside the sample. Furthermore, the inner fourth cube in the agglomerate doesn’t come from a ghosting artifact of the cube in the other frame: their orientation is, indeed, slightly different.
The results in Fig. 2 and Fig. 3 discussed so far demonstrate that Dichography can deal with markedly different shapes and sizes between the two retrieved frames. An additional relevant feature of single-particle CDI experiments at XFELs are the fluctuations in the brightness of the diffraction signals. In fact, samples are intercepting the light pulse while free-flying, and their position in the focus at the diffraction event is ruled by statistics. Particles can intercept the light pulse at varied positions in the focal volume, with distinct light intensity and, consequently, diffraction brightness. This is further enhanced by intensity fluctuations in the XFEL source.
It is straightforward to deduce that, in the “double-hit” case, two particles far apart in the interaction region can be subjected to different pulse intensities, causing both deviations in the overall brightness as well as unbalance in the scattering contribution between the two frames. Less intuitively, this is also a key feature in the two-color case, due to the different positions of the foci of the two pulses, as discussed in the next section (see Sec. D.2 for technical details). These factors must be considered when evaluating the imaging performance of Dichography.
@@
The sensitivity of the method to an imbalance of scattered photons between the two frames, while partially related to the morphological properties of the samples, strongly depends on the overall brightness of the recorded signal and arises from the dependence of noise on the field intensity. In general, a higher photon yield in the signal allows for dichographic reconstructions with a stronger intensity imbalance between the two frames. Further discussion, supported by tests on simulated data, is provided in Sec. S1 of the Supplemental Material [34].
@@
The results on “double-hits” discussed so far, while being of limited scientific interest, represent a strong demonstration of the viability of Dichography for significantly differing morphologies and sample properties. These data is characterized by optimal brightness conditions, with the XFEL operating around its maximum performance in terms of total power per pulse and intensity in the focus. These results represent a solid ground for the discussion of the following dataset, which has a high scientific relevance but poses major challenges due to the advanced operational mode of the XFEL.
III.2 The ultrafast movie of xenon-doped helium nanodroplets
Figure 4: Reconstructions from two-color diffraction patterns acquired at the European XFEL, produced by superfluid helium nanodroplets doped with xenon. Two experimental diffraction patterns are reported in a) and d), on which the diffraction signals produced by two different photon energies, $1.0\text{\,}\mathrm{keV}$ and $1.2\text{\,}\mathrm{keV}$, are superimposed. The intensity is encoded as reported by the color bar at the bottom. The intensity values are translated into numbers of $1.2\text{\,}\mathrm{keV}$ photons per pixel. The inset plot is a zoomed-in region of the diffraction image, to better appreciate the photon statistics, and thus the statistical noise, that affects the experimental data. The two reconstructed frames of the electronic density are reported in b), c) and e), f) for the patterns in a) and d), respectively. In particular, b) and e) report the reconstruction from the $1.0\text{\,}\mathrm{keV}$ pulse, while c) and f) correspond to the $1.2\text{\,}\mathrm{keV}$ pulse, hitting the target $750\text{\,}\mathrm{fs}$ later. The reconstructed density of the xenon doping follows the color bar in Fig. 3. The shape of the helium droplets, independently retrieved before the imaging process and constrained using the DCDI method, are superimposed in blue color.
The experiment was performed at the European XFEL facility, taking advantage of the recently available capability of producing two collinear X-ray pulses of different radiation wavelengths with controllable time delays delay [19]. The two co-propagating pulses are tuned to $1.0\text{\,}\mathrm{keV}$ and $1.2\text{\,}\mathrm{keV}$ photon energy, corresponding to $1.2\text{\,}\mathrm{nm}$ and $1.0\text{\,}\mathrm{nm}$ wavelength, and are separated in time by $750\text{\,}\mathrm{fs}$. Both are focused in the interaction region, where they intercept isolated superfluid helium nanodroplets doped with xenon. An individual droplet in the focus interacts with both pulses, and the corresponding scattered fields incoherently overlap on the detector. Further details on the experiment are discussed in Appendix D.2.
Two diffraction images acquired in these conditions are shown in Fig. 4a and Fig. 4d. A relevant feature of these patterns is the low number of photons recorded by the detector, if compared to the photon yield for the standard, single pulse, operation mode of the XFEL [35] (see Appendix D.2). This aspect represents a challenge for a photon-hungry technique such as CDI, and thus even more so for Dichography (as further investigated in Sec. S1 of the Supplemental Material  [34]).
Nevertheless, the specific properties of the physical system under study, i.e. doped superfluid helium nanodroplets, allow us to take advantage of the Droplet Coherent Diffraction Imaging (DCDI) approach [36]. The DCDI method constrains the underlying helium droplet structure during the actual reconstruction process. The droplet size and brightness are extracted in the pre-processing phase, by fitting the experimental data with the diffraction profile of a spherical sample. The information to be retrieved, therefore, pertains solely to the internal structure of the doping material. For conventional CDI, this greatly reduces the complexity of the problem, it increases the convergence properties of the algorithms compared to standard CDI, and it makes the reconstruction process much more resilient to noise [36].
The DCDI method has been incorporated into the Dichography reconstruction process through a straightforward adaptation of its original workflow, as discussed in Appendix E, and is used here to analyze the two-color diffraction images presented in this section. Its applicability to this dataset relies on the possibility of extracting the droplet size as well as on the persistence of the pristine helium droplet density, embedding the xenon doping, up to a time delay of $750\text{\,}\mathrm{fs}$. Both aspects are investigated and verified using pure-helium data acquired during the same experimental campaign, as reported in Ref. [35].
Despite the demanding experimental conditions, the reduction in problem complexity achieved through the use of DCDI makes these two-color data barely tractable with Dichography in a few optimal cases. The two frames reconstructed from the diffraction pattern shown in Fig. 4a are presented in Fig. 4b for the $1.0\text{\,}\mathrm{keV}$ pulse, and in Fig. 4c for the following $1.2\text{\,}\mathrm{keV}$ pulse. The $1.0\text{\,}\mathrm{keV}$ reconstruction in Fig. 4b reveals a continuous structure extending from the center of the helium droplet to its lower boundary. This feature is consistent with previous observations of xenon atoms aggregating along quantum vortices in superfluid helium [37, 36, 38].
A similar structure is also observed in the $1.2\text{\,}\mathrm{keV}$ reconstruction shown in Fig. 4c, indicating consistency between the two measurements. However, beyond this replicated feature, notable differences are present. In particular, the $1.0\text{\,}\mathrm{keV}$ reconstruction exhibits a high-density spot on the left side of the droplet, which is absent in the $1.2\text{\,}\mathrm{keV}$ reconstruction. Conversely, the $1.2\text{\,}\mathrm{keV}$ reconstruction displays a distinct high-density spot at the lower-right edge of the droplet, with no corresponding feature in the $1.0\text{\,}\mathrm{keV}$ image.
Analogous observations are obtained from the reconstructions obtained from the pattern in Fig. 4d. In particular, here all the high-density regions due to the presence of xenon retrieved in the $1.2\text{\,}\mathrm{keV}$ reconstruction in Fig. 4f can be linked with corresponding features in the $1.0\text{\,}\mathrm{keV}$ reconstruction in 4e. In the earlier $1.0\text{\,}\mathrm{keV}$ frame, however, additional features are visible due to an overall lower signal-to-noise ratio.
[…]
From the representation of the retrieved densities in Fig. 4, the reader may question whether the similar structures in the two frames arise merely as a ghosting artifact – i.e., features from one frame being mistakenly replicated in the other. However, the use of different radiation wavelengths in two-color diffraction data intrinsically implies that the sample is imaged with different spatial resolutions in the two frames. Specifically, the spatial extent corresponding to a single pixel varies between the frames, becoming smaller at higher photon energies due to the increased resolution at a shorter wavelength. As a result, the same structure spans a different number of pixels in the two reconstructions, and appears visually similar in size only when both are displayed with a common scale bar, as done in Fig. 4. This point is further illustrated in Fig. 7 and discussed in more detail in Appendix E. Ghosting in two-color Dichography would then give rise to a replicated structure but with a different scaling. Thus, its presence can be excluded for the xenon structures observed in Fig. 4.
The presence of the same structure in both frames, apart from the artifacts discussed above, implies that the xenon agglomerates survive for at least $750\text{\,}\mathrm{fs}$ after irradiation with the pump pulse. This observation is, however, limited by the achieved spatial resolution of about $20\text{\,}\mathrm{nm}$, which is set by the maximum momentum transfer at which the scattering signal remains above the noise level [35].
IV Discussion
@@
In this work, we introduce the Dichography method to perform Coherent Diffraction Imaging (CDI) in cases where two non-interfering diffraction signals overlap on the detector. This method finds direct application in two-color experiments at XFELs, where pairs of ultrashort pulses with controllable time delay and different wavelengths can be generated. Dichography enables the reconstruction of two independent frames, corresponding to the two distinct views of the system provided by the different colors—thus unlocking the possibility of producing a full-fledged movie of the sample.
@@
The applicability of Dichography extends beyond the two-color case, as it generally applies to any scenario where the recorded diffraction data is the incoherent sum of two independent signals. The so-called “double-hit” case, where two particles, spatially separated in the interaction region, are simultaneously illuminated by a single light pulse, is one such scenario and serves as a benchmark for the method.
@@
In this case, Dichography enables the separation of the scattering contributions and the retrieval of the corresponding densities of the two particles within the focal volume in two separate frames. The various provided examples (see Fig. 2 and Fig. 3) cover a wide range of cases in terms of sample shape, spatial extent and brightness of the two scattering signals. They serve as practical demonstrations that, in principle, no prior knowledge about the samples is required, nor are any specific constraints imposed on the relationship between the properties of the two retrieved frames.
The only assumption about the samples is the oversampling condition, similarly to conventional single-particle CDI – namely, the isolated nature of the nanoparticle in a zero-scattering background. Mathematical considerations suggest that the necessary extent of the surrounding zero-scattering region, known as oversampling degree, is more restrictive for Dichography than for conventional CDI, as further investigated in Appendix B. While the existence of a unique reconstruction is demonstrated for conventional CDI, it is supported by practical evidence in the case of Dichography, and it is yet to be proven. This aspect is discussed in details in Appendix B.1.
@@
The brightness of the patterns and the distribution of the photons between the two frames become significantly more restrictive for the two-color imaging case, were the two overlapping fields come from two XFEL pulses at different photon energies. Here, the special mode of operation of the XFEL undulators delivers two light pulses with significantly reduced intensity with respect to standard operation mode [19], as further discussed in Appendix D.2. Therefore, the reconstructions of xenon-doped superfluid helium nanodroplets shown in Fig. 4 are only possible by employing the DCDI method [36], which is restricted to samples embedded in nanoparticles with known shape and low scattering cross-section – so far only demonstrated for the specific case of superfluid helium nanodroplets [39]. The employment of the DCDI approach greatly reduces the problem complexity and makes the reconstruction process significantly more resilient to noise. Even so, only specific patterns with optimal condition, i.e. higher brightness and similar photon yields between the two frames, turn out to be tractable. Tests on simulated two-color diffraction images, discussed in Sec. S1.4 of the Supplemental Material [34], suggest, however, that even just a twofold increase of the diffraction intensity can lead to a stunning improvement in the reconstruction quality and reliability.
@@
Dichography, similar to conventional single-particle CDI, is an indirect imaging technique, as the sample image has to be retrieved algorithmically from the measured data [6, 7]. The algorithmic approach at the core of Dichography is built on a direct, relatively straightforward, adaptation of the algorithms for single-shot CDI, as discussed in Appendix A. The higher complexity of the dichographic problem and its peculiarities render the algorithm less effective in the reconstruction process. Therefore, the imaging performance is boosted by the combined use with an evolutionary algorithm, which already demonstrated superior performance and reliability for conventional single-particle CDI [40], as further discussed in Appendix C. A more rigorous and fundamental understanding of the problem, which substantially differs from conventional CDI in some key aspects, could result in the design of more efficient imaging algorithms and analysis workflows, extending the range of applicability even further.
V Outlook
@@
Another modern operational mode of XFELs allows for the generation of light pulses with sub-femtosecond durations [5, 4, 33, 3, 2, 1, 41]. The combination of these attosecond pulses with diffraction imaging has already demonstrated the ability to access ultrafast ionization dynamics by exploiting transient electronic resonances [42]. Recently, attosecond pulses have been combined with the two-color mode [1]. Thanks to the short time duration of the pulses, which can outrun the lifetime of excited electronic configurations, two-color attosecond Dichography may provide unprecedented insights into such elusive states, like the lifetimes, spatial distribution and evolution of electronic excitations [43, 44] or plasmonic phenomena such as charge density waves in complex nanostructures, depending on their size, shape and surrounding environment.
@@
Two-color X-ray Dichography can be effectively combined with the use of a third pulse from an optical laser to enable a completely new class of imaging experiments. The laser can be employed to trigger a metastable state, which temporarily opens a specific transition triggered with the XFEL pump and tracked by the XFEL probe. The two consecutive XFEL pulses can also be employed as a double probe. For example, nanoparticles can be superheated by a laser, leading to fast out-of-equilibrium dynamics, like shock-waves and disintegration of the sample [18]. The laser irradiation is followed by the two consecutive XFEL pulses. Dichographic reconstructions are thus two snapshots of the evolving system that allow to precisely track its kinematics and the dynamics of the molecules.
@@
While the current performance of two-color pulses turns out to be challenging for Dichography, recent works have demonstrated the capability of partially separating individual photons between the two pulses. This is achieved by exploiting the sensitivity of the detector to the photon’s energy combined with advanced analysis techniques [35]. The feeding of this additional information into the algorithmic workflow of Dichography is currently under development. Enhancements in the quality and reliability of the reconstructions are expected and they will be the subject of future work.
@@
Improvements of XFELs in two-color mode, in terms of pulse brightness, are also expected in the next years, thanks to the developments of XFEL technology encouraged by the increasing interest of the scientific community [45]. However, even with the current performance, the brightness of the scattering signal can be boosted in different ways. The xenon-doped helium nanodroplets analyzed in this work have an intrinsically low scattering cross section [46]: the highly-scattering xenon represents only a small fraction of the whole system, thus contributing only marginally to the scattered field. An overall diffraction signal even orders of magnitude higher is expected for systems with significantly higher cross sections, e.g. heavier elements. Furthermore, markedly larger scattering cross sections even for lighter materials can be attained at lower photon energies, from few tens to few hundreds of eV [47, 48]. This energy regime in-between the extreme ultraviolet and soft X-ray is also characterized by the presence of electronic resonances for different elements [49], like the biologically relevant carbon and oxygen, where ultrafast electron dynamics can be triggered [42] and traced thanks to the strong variations in the scattering cross section [50].
@@
A further experimental domain that can benefit from the imaging capabilities of Dichography is the study of physical systems sensitive to light polarization. Indeed, several XFEL light sources can partially [51] or fully [52] tune and control the polarization properties of the light in both the XUV [53] and X-ray spectra [54], even with ultrashort light pulses [55]. Moreover, two-color pulses with controllable time-delays and independent polarizations have been recently demonstrated [24]. Dichography can be combined with these experimental capabilities to study the spatial distribution and time evolution of dichroic systems – i.e., systems whose scattering response is sensitive to light polarization. Among these, magnetic domains [56], molecular chirality [57], and orbital ordering in crystals [58] are highly active research topics.
@@
While the difference in wavelength between the two pulses is effective for a better spatial separation of the signal on the scattering detector, it is not a fundamental requirement for Dichography. Thus, if significant structural modifications are expected to take place in the sample between the two XFEL pulses [31], the same photon energy can be potentially employed. This feature would allow the whole set of XFEL undulators to be tuned to the same wavelength, and thus the full peak power to be reached. However the development of methods to produce pairs of electron bunches with sub-picosecond controllable time delay is required to gain full advantage from such an experimental scheme.
@@
Dichography represents a major step forward in harnessing the imaging capabilities of two-color light pulses emerging at XFEL facilities. Indeed, the two views on the sample at different points in time embody, in every respect, one of the original promises of X-ray Free Electron Lasers: recording ultrafast movies of nanomatter.

### v3

@@ II.2 Seeing double from a single shot
The last example on “double-hit” data is reported in Fig. 4d, and presents many peculiarities of the samples shown before combined in one case. The two frames of the reconstruction have a significantly different spatial extension, $80\text{\,}\mathrm{nm}$ for the smaller and up to a maximum of $200\text{\,}\mathrm{nm}$ for the larger. The smaller density has a simple cubic architecture, whereas the larger one can be interpreted as an agglomerate of four nanocubes. The shape of three cubes in the agglomerate is well imprinted into the overall outline, while the presence of the fourth cube is only revealed thanks to the successful reconstruction of the density values inside the sample. Furthermore, the inner fourth cube in the agglomerate does not originate from a ghosting artifact of the cube in the other frame: their orientation is, indeed, slightly different.
The results in Fig. 3 and Fig. 4 discussed so far demonstrate that Dichography can deal with markedly different shapes and sizes between the two retrieved frames. An additional relevant feature of single-particle CDI experiments at XFELs are the fluctuations in the brightness of the diffraction signals. In fact, samples intercept the light pulse while free-flying, and their position in the focus at the diffraction event is ruled by statistics. Particles can intercept the light pulse at varied positions in the focal volume, with distinct light intensity and, consequently, diffraction brightness. This is further enhanced by intensity fluctuations in the XFEL source.
It is straightforward to deduce that, in the “double-hit” case, two particles far apart in the interaction region can be subjected to different pulse intensities, causing both deviations in the overall brightness as well as imbalance in the scattering contribution between the two frames. Less intuitively, this is also a key feature in the two-color case, due to the different positions of the foci of the two pulses, as discussed in the next section (see Sec. V.4.2 for technical details). These factors must be considered when evaluating the imaging performance of Dichography.
@@
The results on “double-hits” discussed so far represent a strong demonstration of the viability of Dichography for significantly differing morphologies and sample properties. This dataset is characterized by optimal brightness conditions, with the XFEL operating around its maximum performance in terms of total power per pulse and intensity in the focus.
III Discussion
@@
In this work, we introduce the Dichography method to perform Coherent Diffraction Imaging (CDI) in cases were two non-interfering diffraction signals overlap on the detector. This method finds direct application in two-color experiments at XFELs, where pairs of ultrashort pulses with controllable time delay and different wavelengths can be generated. Dichography enables the reconstruction of two independent frames, corresponding to the two distinct views of the system provided by the different colors, effectively unlocking the possibility of producing a full-fledged movie of the sample, without a-priori knowledge.
@@
The applicability of Dichography extends beyond the two-color case, as it generally applies to any scenario where the recorded diffraction data is the incoherent sum of two independent signals [33]. The so-called “double-hit” case, where two particles, spatially separated in the interaction region, are simultaneously illuminated by a single light pulse, is one such scenario and serves as a benchmark for the method.
@@
The various examples provided (see Fig. 3 and Fig. 4) cover a wide range of cases in terms of sample shape, spatial extent and brightness of the two scattering signals. They serve as practical demonstrations that, in principle, no prior knowledge about the samples is required, nor are any specific constraints imposed on the relationship between the properties of the two retrieved frames. The only assumption about the samples is their isolation which, similar to conventional single-particle CDI, ensures the oversampling condition.
@@
The brightness of the patterns and the distribution of the photons between the two frames become significantly more restrictive for the two-color imaging case, were the two overlapping fields come from two XFEL pulses at different photon energies. Here, the special mode of operation of the XFEL undulators delivers two light pulses with significantly reduced intensity with respect to standard operation mode [19], as further discussed in Sec. V.4.2. Therefore, the reconstructions of xenon-doped superfluid helium nanodroplets shown in Fig. 2 are only possible by employing the DCDI method [40], which is restricted to samples embedded in nanoparticles with known shape and low scattering cross-section – so far only demonstrated for the specific case of superfluid helium nanodroplets [44].
The employment of the DCDI approach greatly reduces the problem complexity and makes the reconstruction process significantly more resilient to noise. Under the experimental conditions of our measurement, only specific patterns with optimal condition, i.e., higher brightness and similar photon yields between the two frames, turned out to be tractable (as further discussed in Sec. S2 of the Supplemental Material [39]).
Tests on simulated two-color diffraction images, however, suggest that even just a twofold increase of the diffraction intensity can lead to a radical improvement in the reconstruction quality and reliability (see Sec. S1.4 of the Supplemental Material [39]).
@@
Dichography, similar to conventional single-particle CDI, is an indirect imaging technique, as the sample image has to be retrieved algorithmically from the measured data [6, 7]. The algorithmic approach at the core of Dichography is built on a direct, relatively straightforward, adaptation of the algorithms for single-shot CDI, as discussed in Sec. V.1. The higher complexity of the dichographic problem and its peculiarities render the algorithm less effective in the reconstruction process. Therefore, the imaging performance is enhanced by coupling with an evolutionary algorithm, which already demonstrated superior performance and reliability for conventional single-particle CDI [45], as further discussed in Sec. V.3. A more rigorous and fundamental understanding of the problem, which substantially differs from conventional CDI in some key aspects, could result in the design of more efficient imaging algorithms and analysis workflows, extending the range of applicability even further.
IV Outlook
@@
Another modern operational mode of XFELs allows for the generation of light pulses with sub-femtosecond durations [5, 4, 43, 3, 2, 1, 46]. The combination of these attosecond pulses with diffraction imaging has already demonstrated the ability to access ultrafast ionization dynamics by exploiting transient electronic resonances [47]. Recently, attosecond pulses have been combined with the two-color mode [1]. Thanks to the short time duration of the pulses, which can outrun the lifetime of excited electronic configurations, two-color attosecond Dichography may provide unprecedented insights into such elusive states, including the possibility to trace the spatial distribution and temporal evolution of electronic excitations [48, 49] or plasmonic phenomena such as charge density waves in complex nanostructures, depending on their size, shape and surrounding environment.
@@
Two-color X-ray Dichography can be effectively combined with the use of a third pulse from an optical laser to enable a completely new class of imaging experiments. The laser can be employed to trigger a metastable state, which temporarily opens a specific transition triggered with the XFEL pump and tracked by the XFEL probe. The two consecutive XFEL pulses can then be employed as a double probe. For example, nanoparticles can be superheated by an intense optical laser pulse, leading to fast out-of-equilibrium dynamics, like shock-waves and disintegration of the sample [18]. The XFEL pulses can either interact with the sample before and after, or both following, or even overlapping, laser irradiation. Dichographic reconstructions are thus two snapshots of the evolving system that allow for the precise tracking its kinematics and the dynamics of the molecules.
@@
While the current performance of two-color pulses turns out to be challenging for Dichography, recent works have demonstrated the capability of partially separating individual photons between the two pulses directly in the recorded signal [38]. This is achieved by exploiting the sensitivity of the detector to the photon’s energy combined with advanced analysis techniques, and can be employed only in the regions of the scattering signal where photon events are sufficiently rare [38], typically at high scattering angles. Integrating this additional information into the algorithmic workflow of Dichography is currently under development. Enhancements in the quality and reliability of the reconstructions are foreseen and they will be the subject of future work.
@@
Improvements of XFELs in two-color mode, in terms of pulse brightness, are also expected in the next years, thanks to the developments of XFEL technology encouraged by the increasing interest of the scientific community [50]. However, even with the current performance, the brightness of the scattering signal can be boosted in different ways. The xenon-doped helium nanodroplets analyzed in this work have an intrinsically low scattering cross section [51]: the highly-scattering xenon represents only a small fraction of the whole system, thus contributing only marginally to the scattered field. An overall diffraction signal even orders of magnitude higher is expected for systems with significantly higher cross sections, e.g. heavier elements. Furthermore, markedly larger scattering cross sections even for lighter materials can be attained at lower photon energies, from a few tens to few hundreds of eV [52, 53]. This energy regime in-between the extreme ultraviolet and soft X-ray is also characterized by the presence of electronic resonances for different elements [54], like the biologically relevant carbon and oxygen, where ultrafast electron dynamics can be triggered [47] and traced thanks to the strong variations in the scattering cross section [55].
@@
A further experimental domain that can benefit from the imaging capabilities of Dichography is the study of physical systems sensitive to light polarization. Indeed, several XFEL light sources can partially [56] or fully [57] tune and control the polarization properties of the light in both the XUV [58] and X-ray spectra [59], even with ultrashort light pulses [60]. Moreover, two-color pulses with controllable time-delays and independent polarizations have been recently demonstrated [24]. Dichography can be combined with these experimental capabilities to study the spatial distribution and time evolution of dichroic systems, i.e., systems whose scattering response is sensitive to light polarization. Among these, magnetic domains [61], molecular chirality [62], and orbital ordering in crystals [63] are highly active research topics.
@@
While the difference in wavelength between the two pulses is effective for a better spatial separation of the signal on the scattering detector, it is not a fundamental requirement for Dichography [33]. In particular, if significant structural modifications are expected to take place in the sample between the two XFEL pulses [31], the same photon energy can be effectively employed. This feature would allow the whole set of XFEL undulators to be tuned to the same wavelength, and thus the full peak power to be reached. However, the development of methods to produce pairs of electron bunches with sub-picosecond controllable time delay is required to gain full advantage from such an experimental scheme.
Dichography can be in principle extended to 3D imaging, for which the imaging problem is more constrained, i.e., easier to solve [37]. Furthermore, it can be extended to the reconstruction of more than two frames [33], especially for sample properties characterized by a high constraint ratio [35] (like doped helium nanodroplets, see Sec. V.2.2), exploiting future, more advanced, multicolor capabilities of modern XFELs, like the possibility of delivering more than two different photon energies in the interaction region.
@@
Dichography represents a major step forward in harnessing the imaging capabilities of two-color light pulses emerging at XFEL facilities, enabling a new class of time-resolved imaging experiments. Indeed, the two views on the sample at different points in time embody, in every respect, one of the original promises of X-ray Free Electron Lasers: recording ultrafast movies of nanomatter.